# suppReferences

### v3

@@ Discussion
While concerns have arisen about LLMs in facilitating the creation of misinformation[42, 43], our work demonstrates LLMs’ potential to improve the online information ecosystem by correcting misinformation[1, 2]. Real-world misinformation can combine (in)accurate, factually accurate but misleading, and unverifiable claims. Identifying a piece of content as misinformation without further interventions, done by many previous AI-driven approaches[44], has limited impact on reducing misinformation’s detrimental effects. Showing the identification results without explanations might even increase false beliefs[45]. Correcting misinformation has been shown to reduce its spread and false beliefs, whose effectiveness is affected by quality, timeliness, and scalability[4, 22]. However, research on scaling up and accelerating misinformation correction—which may include automation—is still at the early stage. Existing AI models struggle to identify and correct misinformation, especially on social media, where misinformation is not restricted to narrow domains and spreads rapidly. Correcting misinformation requires comprehending content that can be multimodal and the context beyond it that often involves emerging events. It requires identifying what part(s) of the content is (in)accurate and explaining why that part of the content is (in)accurate with trustworthy references. We propose Muse and demonstrate the high quality of its automatically generated responses to social media posts that potentially are misinformation. Muse augments existing powerful LLMs (here, it is GPT-4) with the capabilities of addressing images, accessing up-to-date knowledge, and finding accurate references. Results further validate that GPT-4 struggles to effectively respond to visual content but Muse excels (Fig. 3a). Muse also exhibits significantly fewer “hallucinations” by having fewer errors in identifying and explaining (in)accuracies (Fig. 2c,e), generating text that is more factual and relevant to the responded content (Fig. 2g-h), and providing more references that are real and relevant to the generated text (Fig. 2l-m) than GPT-4.
@@
We provide Muse as a solution to assist social media users and platforms in accurately, scalably, promptly, and transparently responding to suspicious content. Muse is end-to-end and thereby simple to use. It is nonparametric, i.e., does not need training or fine-tuning an AI model, and thereby easily and cheaply updated, especially compared to parametric models[32]. Meanwhile, our results reveal that Muse-generated responses have high quality in identifying and explaining inaccuracies, generated text, and provided references, significantly surpassing GPT-4 and laypeople who produce even high-helpfulness responses in correcting misinformation across modalities and domains (Fig. 2-3). Besides the highest accuracy and factuality (Fig. 2c,e,g-h,l-m), Muse’s generated responses show the highest readability by being the most explicit, fluent, and coherent (Fig. 2b,i-j). Muse can reduce the risk of a correction backfiring (i.e., reinforcing rather than reducing people’s false beliefs) by generating responses with the least toxic text and most credible references[28, 38, 39] (Fig. 2k,n). It can also reduce the implied truth effect (i.e., increasing people’s perceived accuracy of overlooked inaccuracies)[22] by comprehensively identifying all the inaccuracies in a social media post (Fig. 2d) and being more capable of correcting misinformation at scale. These high-quality responses by Muse can generally be obtained two minutes after suspicious content (Methods). By transparently providing references that refer to the retrieved web pages where evidence was collected, users can become more informed and also verify responses themselves (Fig 2,3f).
Muse’s responses cost about 0.5 USD per social media post at the time of our evaluation, though this cost has now been reduced to 0.2 USD, as GPT-4’s price has lowered (Methods). Our focus in designing Muse was in maximizing the quality of corrections. Considering the task’s complexity and our significant improvement in quality, the cost is relatively inexpensive compared to alternatives. For example, a crowd of laypeople can already cost about 0.9 USD[19] to identify whether a new article’s headline and lede contains misinformation without writing down the explanation.
@@
This work also faces the following limitations. First, although Muse is capable of responding to multimodal misinformation with text and images, it cannot accept video inputs. Second, we only focus on English, one of the most spoken languages in the world. Third, we evaluated Muse using real social media content on a single platform, X, as its Community Notes system has been shown to reduce the spread of misinformation[17] and transparent. X is also a popular social media platform, and one where more than half of users consume news regularly[46] and where misinformation has been shown to diffuse faster than the truth[13]. Fourth, experts assessed and compared Muse against one other LLM, GPT-4, which can also be seen as an ablation study, since Muse augments GPT-4 (Methods). GPT-4 was chosen as a comparison as it is one of the best performing LLMs across a wide range of tasks available today[47, 48].
@@ Conclusion
We proposed Muse, a nonparametric LLM to scale up and accelerate misinformation corrections on social media. Muse can integrate textual and visual information, access timely knowledge, and generate responses that use natural languages and have accurate and trustworthy references. Experts assessed that Muse significantly outperforms GPT-4 and even laypeople who produce high-helpfulness responses on X Community Notes in identifying and explaining inaccuracies, generating high-quality text, providing high-quality references, and producing the highest quality of responses overall. Muse also excels when the responded content is textual or multimodal (textual and visual) and when it is related to politics and international affairs, economy and business, crime and law, social issues and human rights, or health and medicine. This study demonstrates the potential of LLMs in responding to online misinformation effectively, scalably, promptly, and transparently.
@@ Acknowledgements
(a) Quality of responses in identifying and explaining (in)accuracies.
(b) Quality of responses in generated text.
(c) Quality of responses in references.
Figure S7: Expert evaluation results for textual posts (#=155).
(a) Quality of responses in identifying and explaining (in)accuracies.
(b) Quality of responses in generated text.
(c) Quality of responses in references.
Figure S8: Expert evaluation results for multimodal (textual and visual) posts (#=77).
(a) Quality of responses in identifying and explaining (in)accuracies.
(b) Quality of responses in generated text.
(c) Quality of responses in references.
Figure S9: Expert evaluation results for political and international-affair posts (#=80).
(a) Quality of responses in identifying and explaining (in)accuracies.
(b) Quality of responses in generated text.
(c) Quality of responses in references.
Figure S10: Expert evaluation results for economic and business posts (#=38).
(a) Quality of responses in identifying and explaining (in)accuracies.
(b) Quality of responses in generated text.
(c) Quality of responses in references.
Figure S11: Expert evaluation results for crime and law posts (#=38).
(a) Quality of responses in identifying and explaining (in)accuracies.
(b) Quality of responses in generated text.
(c) Quality of responses in references.
Figure S12: Expert evaluation results for social-issue and human-right posts (#=30).
(a) Quality of responses in identifying and explaining (in)accuracies.
(b) Quality of responses in generated text.
(c) Quality of responses in references.
Figure S13: Expert evaluation results for health and medicine posts (#=24).
@@
1Describe an image in an informative way. Your description should be only based on the given {short caption}, {name of each person}, and {raw text}. If the image is from social media, you should start with "A screenshot of". If the image is a quote from someone, you should start with "A quote from" followed by this person’s name if there is any, then by the quoted text. If the image is an article, you should start with "An article". If the image is a photo, you should start with "A photo of". If the image is a map, you should start with "A map of". {raw text} may contain nonsense data that are unnecessarily included in the image description; however, {name of each person} is not, and if the concept in {raw text} has a conflict with that in {short caption} (e.g., "Robbie Lemos" versus "robbie leems" shown later), {raw text} is often the right one.2short caption: {a woman with glasses and a quote that says, in real life, i assure you there is no such thing as algebra}3name of each person: {Fran Lebowitz}4raw text: {"In real life, I assure you, there is no such thing as algebra."}5image description: {A quote from Fran Lebowitz, "In real life, I assure you, there is no such thing as algebra."}6short caption: {two men in suits}7name of each person: {Jim Caviezel, Michael Emerson}8raw text: {}9image description: {A photo of Jim Caviezel and Michael Emerson in suits}10short caption: {robbie leems on twitter}11name of each person: {}12raw text: {Robbie Lemos @RobbieLemos 1d I’d like to congratulate my dear friend Deep Mind on a wonderful 1st day at work today at Google. Just in time for #EarthDay2023, cheers brother! 1 2 3,790}13image description: {A screenshot of a post of Robbie Lemos, "I’d like to congratulate my dear friend Deep Mind on a wonderful 1st day at work today at Google. Just in time for #EarthDay2023, cheers brother!" The post was posted on Twitter.}14short caption: {a moose}15name of each person: {}16raw text: {Yahoo Finance @YahooFinance Typically, the stock market bottoms four to five months before a recession ends, but RBC’s research details that it has bottomed as early as nine months before the end of a recession. finance.yahoo.com Could the stock market power through a recession? ’This would be rare.’ 09:57 22/4/2023 3.4,011 Views 1 Retweet 1 Quote 5 Likes 1 Bookmark}17image description: {A screenshot of a post from Yahoo Finance, "Typically, the stock market bottoms four to five months before a recession ends, but RBC’s research details that it has bottomed as early as nine months before the end of a recession." The post shared an article from finance.yahoo.com claiming, "Could the stock market power through a recession? ’This would be rare.’" with a picture of a moose. The post was posted at 09:57 22/4/2023.}18short caption: {a person pouring tea into a cup}19name of each person: {}20raw text: {New research reveals how coffee and tea can affect risk of early death for adults with diabetes By Sandee LaMotte, CNN Updated 7:01 PM EDT, Wed April 19, 2023 f The health benefits of tea 01:10 - Source: CNN}21image description: {An article claiming, "New research reveals how coffee and tea can affect risk of early death for adults with diabetes." It attached a picture of a person pouring tea into a cup. It was written by Sandee LaMotte, published by CNN, and updated at 7:01 PM EDT, Wed April 19, 2023.}22short caption: {two people standing next to each other with the words love is blind}23name of each person: {Nick Lachey}24raw text: {\"Love Is Blind\" co-host faceplants with a regressive line of questioning Hayley Miller MSNBC DAILY MSNBC}25image description: {An article claiming, "’Love Is Blind’ co-host faceplants with a regressive line of questioning." It attached a picture of Nick Lachey and another person standing next to each other. It was written by Hayley Miller and published by MSNBC.}26short caption: {a bar graph that shows how engaged are the most followed journalists on twitter}27name of each person: {Rahul Kanwal}28raw text: {How engaged are the most-followed journalists on Twitter? Percentage of tweets from each journalist that are at-replies BDUTT 64% RealMikeWilbon 38% 34% ErinAndrews 31% stephenasmith 24% andersoncooper rahulkanwal 13% 13% ninagarcia 5% sardesairajdeep 1% BillSimmons maddow 1% 0% 10% 20% 30% 40% 50% 60% chart by mattmaldre.com}29image description: {A bar graph showing how engaged the most followed journalists, including Rahul Kanwal, are on Twitter through the percentage of tweets from each journalist that are at-replies. The chart was made by mattmaldre.com.}30short caption: {a graph showing the global defense budget by region}31name of each person: {}32raw text: {Global Defense Budgets by Region (\$ Billions) \$1,000 800 600 400 200 0 2020 2021 2022 2023 2024 2025 Asia-Pacific Latin America North America Sub-Saharan Africa Europe Middle East & North Africa Russia & Commonwealth of Independent States Source: Aviation Week}33image description: {A graph showing the global defense budget by region. It is from Aviation Week.}34short caption: {[IMAGE_CAPTION]}35name of each person: {[CELEBRITIES]}36raw text: {[OCR]}37image description:
@@
Figure S27: Distribution of post times of tweets in X Community Notes included in our analyses.

### v4

@@ Discussion
While concerns have arisen about LLMs in facilitating the creation of misinformation[48, 49], our work demonstrates LLMs’ potential to improve the online information ecosystem by correcting misinformation[2, 3]. Correcting misinformation requires identifying what part(s) of the content is (in)accurate and explaining why that part of the content (in)accurate with trustworthy references. Such identifications with clear explanations have been shown to reduce misinformation’s spread and people’s false beliefs, but its effectiveness is affected by quality, timeliness, and scalability[4, 22]. However, research on scaling up and accelerating misinformation correction is still at the early stage. Existing AI models are ill-equipped to identify and correct misinformation on social media. Such misinformation often spreads rapidly and is not restricted to narrow domains. It can combine (in)accurate, factually accurate but misleading, and unverifiable claims. It can employ various tactics, deliberately or accidentally, to create falsehood or misleadingness. Addressing it requires comprehending content that can be multimodal and the context beyond it that often involves emerging events. Existing methods to evaluate the quality of corrections presented in natural language also struggle to be accurate and comprehensive. Such evaluation relies on professional knowledge and skills. It should thoroughly assess the identification and explanation of (in)accuracies, as well as the use of references from diverse perspectives.
We propose Muse, which augments existing powerful LLMs (here, it is GPT-4) with the capabilities of addressing images, accessing up-to-date knowledge, and finding accurate references, and 13 dimensions of misinformation correction quality. Evaluation by fact-checking experts demonstrates the high quality of Muse’s automatically generated responses to social media posts that potentially are misinformation. Results further validate that GPT-4 struggles to effectively respond to visual content but Muse excels (Fig. 3a). Muse also exhibits significantly fewer “hallucinations” by having fewer errors in identifying and explaining (in)accuracies (Fig. 2c,e), generating text that is more factual and relevant to the responded content (Fig. 2g-h), and providing more references that are real and relevant to the generated text (Fig. 2l-m) than GPT-4. Note that addressing LLM ’hallucination’ has been a long-standing challenge stressed in many works[26, 50].
@@
We provide Muse as a solution to assist social media users and platforms in reliably, scalably, promptly, and transparently responding to suspicious content. Muse is end-to-end and thereby simple to use. It is nonparametric and thereby easily and cheaply updated[51]. Meanwhile, our results reveal that Muse-generated responses have high quality in identifying and explaining inaccuracies, generated text, and provided references, significantly surpassing GPT-4 and laypeople who produce even high-helpfulness responses in correcting misinformation across modalities, domains, political leanings, and tactics and that has not even been fact-checked online (Fig. 2-3). Besides the highest accuracy and factuality (Fig. 2c,e,g-h,l-m), Muse’s generated responses show the highest readability by being the most explicit, fluent, and coherent (Fig. 2b,i-j). Muse can reduce the risk of a correction backfiring (i.e., reinforcing rather than reducing people’s false beliefs) by generating responses with the least toxic text and most credible references[29, 41, 42] (Fig. 2k,n). It can also reduce the implied truth effect (i.e., increasing people’s perceived accuracy of overlooked inaccuracies)[22] by comprehensively identifying all the inaccuracies in a social media post (Fig. 2d) and being more capable of correcting misinformation at scale. These high-quality responses by Muse can generally be obtained within minutes of suspicious content appearing on social media (Methods). By transparently providing references that refer to the retrieved web pages where evidence was collected, users can become more informed and also verify responses themselves (Fig 2,3f). Muse can also inform users of the current lack of evidence, provide an accuracy nudge, and when applicable, express uncertainty (Supp. Fig. S8).
Muse’s responses cost about 0.5 USD per social media post at the time of our evaluation, though this cost has now been reduced to 0.2 USD, as GPT-4’s price has lowered (as of February 2024; Methods). Our focus in designing Muse was in maximizing the quality of corrections. Considering the task’s complexity and our significant improvement in quality, the cost is relatively inexpensive compared to alternatives. For example, a crowd of laypeople can already cost about 0.9 USD[19] to identify whether a new article’s headline and lede contains misinformation without writing down the explanation. Notably, we also changed Muse’s foundation LLM from GPT-4 to open-sourced Llama-3 (70B) and observed two Muses are comparable without modifying any other implementation details (Supp. Fig. S9; Methods), suggesting Muse’s generalizability. Note that the knowledge cutoff for Llama-3 is December 2023, whereas that of GPT-4 is September 2021. Llama-3’s training data might have included some or even all tweets used in evaluation and related events, knowledge, and corrections (e.g., their Community Notes data). As a comparison, GPT-4’s training data might only include up to 10% (Methods, where we show that Muse’s performance remains consistent after removing these tweets). Hence, it is unfair to compare between Llama-3-based and GPT-4-based Muse, and we did not include Llama-3 and its based Muse in expert evaluation. Nevertheless, our qualitative analysis illustrated in Supp. Fig. S9 reveals the potential to significantly reduce Muse’s cost, as GPT-4 was Muse’s most costly component. Furthermore, as one of the most capable open-sourced LLMs, Llama-3 still cannot accept visual inputs (as of August 2024), emphasizing the importance of Muse’s image describer module.
@@
This work also faces the following limitations. First, although Muse is capable of responding to multimodal misinformation with text and images, it cannot accept video inputs. Second, we only focus on English, one of the most spoken languages in the world. Third, we evaluated Muse using real social media content on a single platform, X, as its Community Notes system has been shown to reduce the spread of misinformation[17] and transparent. X is also a popular social media platform, and one where more than half of users consume news regularly[52] and where misinformation has been shown to diffuse faster than the truth[13]. Fourth, experts assessed and compared Muse against one foundation LLM, GPT-4, which can also be seen as an ablation study, since Muse augments GPT-4 (see Methods and Supp. Fig. S24 for additional results of our ablation study). GPT-4 was chosen as a comparison as it is one of the best performing LLMs across a wide range of tasks available today[53, 54].
@@ Conclusion
We proposed Muse, a nonparametric LLM to identify and explain (in)accuracies in social media content. Muse can integrate textual and visual information, access timely knowledge, and generate responses that use natural languages and have accurate and trustworthy references. Experts assessed that Muse significantly outperforms GPT-4 and even laypeople who write high-helpfulness responses on X Community Notes in identifying and explaining (in)accuracies, generating high-quality text, providing high-quality references, and producing high-quality responses overall. Muse excels when the responded content is textual or multimodal, related to a broad range of domains, false or misleading by applying various tactics, and even when the content has not been fact-checked online. Muse also performs consistently well when responding to liberal and conservative content. This work demonstrates the potential of LLMs to respond to online misinformation in a scalable, prompt, reliable, and transparent manner—addressing key bottlenecks where existing studies struggle. This work also contributes to a comprehensive framework, incorporating with 13 associated standards (Methods, where we analyzed their importance), for evaluating misinformation correction models. Additionally, we publicly release rich annotations on social media posts covering aspects such as political leanings, domains, and tactics that make them false or misleading, along with expert-annotated responses to enable research.
@@ Acknowledgements
Figure S7: Examples of social media content that uses various tactics to make it or part of it false or misleading. These tactics include a: Fabricating a news event (#1), a story with an authentic image (#2, also related to conspiracy theory), and an image with AI (#3). b: Lacking context regarding the motivation behind behavior. Here, it is the consideration that the bill contained no funding for Florida, without which could mislead readers. c: Misinterpreting or misrepresenting a plan, where Biden did not promise to cancel all the debt (#1), or symbol, which is not Nazi SS runes but a shorthand for the 46th Separate Airmobile Brigade (#2). d: Using loaded language. e: Improperly analogizing or equating masks blocking respiratory droplets and underwear blocking gas molecules (#1), and COVID-19 and cancer (#2). f: Presenting false, partial, or biased data. 365348 includes multiple reports for the same child and runaways who may not be considered missing because of their guardians know their whereabouts (#1). The data ends in 1885 and is from a specific high elevation site in Greenland, not global temperatures (#2). g: Implying false or oversimplified causation. Many COVID-19 deaths occurred among Medicare beneficiaries because Medicare primarily serves the groups who are at greater risk of adverse outcomes from COVID-19 (#1). The boy was shot when—but not because—he was campaigning for Raphael Warnock (#2). Note that one content may apply more than one tactic intentionally or accidentally.
Figure S8: Examples of Muse-generated responses to social media posts that lack evidence to verify.
Figure S9: Comparison between responses generated by GPT-4, Llama-3, and their based Muse. We observed that GPT4-based Muse and Llama3-based Muse frequently generate similar responses to social media content (see the top example). GPT4-based Muse occasionally generates more comprehensive corrections than Llama3-based Muse (see the middle example). Note that the knowledge cutoff for Llama3 is December 2023, whereas that of GPT-4 is September 2021. In other words, Llama-3’s training data include more recent events and knowledge than GPT-4’s (see the bottom example, where the post was posted in February 2023, and Llama-3 more accurately identifies its falsehood than GPT-4).
@@
1Describe an image in an informative way. Your description should be only based on the given {short caption}, {name of each person}, and {raw text}. If the image is from social media, you should start with "A screenshot of". If the image is a quote from someone, you should start with "A quote from" followed by this person’s name if there is any, then by the quoted text. If the image is an article, you should start with "An article". If the image is a photo, you should start with "A photo of". If the image is a map, you should start with "A map of". {raw text} may contain nonsense data that are unnecessarily included in the image description; however, {name of each person} is not, and if the concept in {raw text} has a conflict with that in {short caption} (e.g., "Robbie Lemos" versus "robbie leems" shown later), {raw text} is often the right one.2short caption: {a woman with glasses and a quote that says, in real life, i assure you there is no such thing as algebra}3name of each person: {Fran Lebowitz}4raw text: {"In real life, I assure you, there is no such thing as algebra."}5image description: {A quote from Fran Lebowitz, "In real life, I assure you, there is no such thing as algebra."}6short caption: {two men in suits}7name of each person: {Jim Caviezel, Michael Emerson}8raw text: {}9image description: {A photo of Jim Caviezel and Michael Emerson in suits}10short caption: {robbie leems on twitter}11name of each person: {}12raw text: {Robbie Lemos @RobbieLemos 1d I’d like to congratulate my dear friend Deep Mind on a wonderful 1st day at work today at Google. Just in time for #EarthDay2023, cheers brother! 1 2 3,790}13image description: {A screenshot of a post of Robbie Lemos, "I’d like to congratulate my dear friend Deep Mind on a wonderful 1st day at work today at Google. Just in time for #EarthDay2023, cheers brother!" The post was posted on Twitter.}14short caption: {a moose}15name of each person: {}16raw text: {Yahoo Finance @YahooFinance Typically, the stock market bottoms four to five months before a recession ends, but RBC’s research details that it has bottomed as early as nine months before the end of a recession. finance.yahoo.com Could the stock market power through a recession? ’This would be rare.’ 09:57 22/4/2023 3.4,011 Views 1 Retweet 1 Quote 5 Likes 1 Bookmark}17image description: {A screenshot of a post from Yahoo Finance, "Typically, the stock market bottoms four to five months before a recession ends, but RBC’s research details that it has bottomed as early as nine months before the end of a recession." The post shared an article from finance.yahoo.com claiming, "Could the stock market power through a recession? ’This would be rare.’" with a picture of a moose. The post was posted at 09:57 22/4/2023.}18short caption: {a person pouring tea into a cup}19name of each person: {}20raw text: {New research reveals how coffee and tea can affect risk of early death for adults with diabetes By Sandee LaMotte, CNN Updated 7:01 PM EDT, Wed April 19, 2023 f The health benefits of tea 01:10 - Source: CNN}21image description: {An article claiming, "New research reveals how coffee and tea can affect risk of early death for adults with diabetes." It attached a picture of a person pouring tea into a cup. It was written by Sandee LaMotte, published by CNN, and updated at 7:01 PM EDT, Wed April 19, 2023.}22short caption: {two people standing next to each other with the words love is blind}23name of each person: {Nick Lachey}24raw text: {\"Love Is Blind\" co-host faceplants with a regressive line of questioning Hayley Miller MSNBC DAILY MSNBC}25image description: {An article claiming, "’Love Is Blind’ co-host faceplants with a regressive line of questioning." It attached a picture of Nick Lachey and another person standing next to each other. It was written by Hayley Miller and published by MSNBC.}26short caption: {a bar graph that shows how engaged are the most followed journalists on twitter}27name of each person: {Rahul Kanwal}28raw text: {How engaged are the most-followed journalists on Twitter? Percentage of tweets from each journalist that are at-replies BDUTT 64%29image description: {A bar graph showing how engaged the most followed journalists, including Rahul Kanwal, are on Twitter through the percentage of tweets from each journalist that are at-replies. The chart was made by mattmaldre.com.}30short caption: {a graph showing the global defense budget by region}31name of each person: {}32raw text: {Global Defense Budgets by Region (\$ Billions) \$1,000 800 600 400 200 0 2020 2021 2022 2023 2024 2025 Asia-Pacific Latin America North America Sub-Saharan Africa Europe Middle East & North Africa Russia & Commonwealth of Independent States Source: Aviation Week}33image description: {A graph showing the global defense budget by region. It is from Aviation Week.}34short caption: {[IMAGE_CAPTION]}35name of each person: {[CELEBRITIES]}36raw text: {[OCR]}37image description:
@@
(a) Correlation for all responses.
(b) Correlation for responses whose overall quality is relatively low vs high.
Figure S25: Spearman correlation coefficient between overall response quality and each criterion that specifies overall response quality.